# Taehoon Yoon     Yunhong Min     Kyeongmin Yeo     Minhyuk Sung KAIST

### v2

@@ Appendix E Additional Evaluation Results of MALA and pCNL Initializations under Varying Step Size
AcceptanceRefer to caption T2I-Count [73]Refer to caption MAE [75]Refer to caption LPIPS MPD [79]Refer to caption
@@
AcceptanceRefer to caption Aesthetic Score [76]Refer to caption LPIPS MPD [79]Refer to caption
@@
Appendix F Ablation Study: Varying NFE Allocation for Initial Particle Sampling Stage
In this section, we present an ablation study examining how performance varies with different allocations of the total number of function evaluations (NFE) between the initial particle sampling stage (via Top-$K$-of-$N$ or MCMC) and the subsequent SMC stage. In our main experiments, we adopt a balanced allocation, with 50% of the NFE budget used for initial particle sampling and the remaining 50% for SMC (denoted as 50%/50%).
To assess the effect of this design choice, we evaluate two alternative NFE splits: 25% for initial particle sampling and 75% for SMC (denoted as 25%/75%), and 75% for initial particle sampling with only 25% for SMC (denoted as 75%/25%). We conduct this analysis on both Top-$K$-of-$N$ and $\Psi$-Sampler. For consistency, we fix the number of SMC steps to 25, adjusting the number of particles accordingly: 30 particles for the 25%/75% setting and 10 particles for the 75%/25% setting. For the aesthetic-preference image generation, we use half the number of particles in both settings to reflect the halved NFE. Tab. 2 summarizes the results.
@@ Appendix F Ablation Study: Varying NFE Allocation for Initial Particle Sampling Stage
[…]
@@
Table 2: Ablation study results on varying NFE allocation conducted with $\Psi$-Sampler and Top-$K$-of-$N$ across three task domains. Bold indicates the best performance, while underline denotes the second-best result for each metric. Metrics marked with are used as seen reward during reward-guided sampling, where others are held-out reward. Higher values indicate better performance ($\uparrow$), unless otherwise noted ($\downarrow$).
Appendix G $\Psi$-Sampler with Other Score-Based Generative Model
@@
[…]
@@
FreeDoM [40] TDS [14] DAS [12] Top-$K$-of-$N$ ULA MALA $\Psi$-Sampler
@@
Appendix H Additional Qualitative Results
@@
Layout FreeDoM [40] TDS [14] DAS [12] Top-$K$-of-$N$ ULA MALA $\Psi$-Sampler
@@
FreeDoM [40] TDS [14] DAS [12] Top-$K$-of-$N$ ULA MALA $\Psi$-Sampler
@@
FreeDoM [40] TDS [14] DAS [12] Top-$K$-of-$N$ ULA MALA $\Psi$-Sampler
@@
FreeDoM [40] TDS [14] DAS [12] Top-$K$-of-$N$ ULA MALA $\Psi$-Sampler

### v3

@@ Appendix F Additional Evaluation Results of MALA and pCNL Initializations under Varying Step Size
Acceptance
T2I-Count [73]
MAE [75]
LPIPS MPD [79]
@@
Acceptance
Aesthetic Score [76]
LPIPS MPD [79]
@@
Appendix G Ablation Study: Varying NFE Allocation for Initial Particle Sampling Stage
In this section, we present an ablation study examining how performance varies with different allocations of the total number of function evaluations (NFE) between the initial particle sampling stage (via Top-$K$-of-NN or MCMC) and the subsequent SMC stage. In our main experiments, we adopt a balanced allocation, with 50% of the NFE budget used for initial particle sampling and the remaining 50% for SMC (denoted as 50%/50%).
To assess the effect of this design choice, we evaluate two alternative NFE splits: 25% for initial particle sampling and 75% for SMC (denoted as 25%/75%), and 75% for initial particle sampling with only 25% for SMC (denoted as 75%/25%). We conduct this analysis on both Top-$K$-of-NN and $\Psi$-Sampler. For consistency, we fix the number of SMC steps to 25, adjusting the number of particles accordingly: 30 particles for the 25%/75% setting and 10 particles for the 75%/25% setting. For the aesthetic-preference image generation, we use half the number of particles in both settings to reflect the halved NFE. Tab. 2 summarizes the results.
@@ Appendix G Ablation Study: Varying NFE Allocation for Initial Particle Sampling Stage
Quantity Aware $\text{T2I-Count}$$^{\dagger}$ [73] $\downarrow$ 1.021 1.077 2.934 0.804 0.850 1.892
@@
Table 2: Ablation study results on varying NFE allocation conducted with $\Psi$-Sampler and Top-$K$-of-NN across three task domains. Bold indicates the best performance, while underline denotes the second-best result for each metric. Metrics marked with are used as seen reward during reward-guided sampling, where others are held-out reward. Higher values indicate better performance ($\uparrow$), unless otherwise noted ($\downarrow$).
Appendix H $\Psi$-Sampler with Other Score-Based Generative Model
@@
Single Particle SMC-Based Methods Tasks Metrics Sampling from Prior Sampling from Posterior DPS [39] FreeDoM [40] TDS [14] DAS [12] Top-$K$-of-NN ULA MALA $\Psi$-Sampler Layout to Image $\text{GroundingDINO}$$^{\dagger}$ [72] $\uparrow$ 0.144 0.159 0.403 0.338 0.406 0.388 0.392 0.429 mIoU [74] $\uparrow$ 0.229 0.242 0.405 0.343 0.406 0.393 0.394 0.432 ImageReward [77] $\uparrow$ 1.241 1.068 1.363 1.263 1.478 1.227 1.326 1.502 VQA [78] $\uparrow$ 0.779 0.754 0.835 0.808 0.851 0.776 0.832 0.853 Quantity Aware $\text{T2I-Count}$$^{\dagger}$ [73] $\downarrow$ 11.290 12.839 0.110 0.122 0.0628 0.220 0.148 0.027 MAE [75] $\downarrow$ 12.3 13.8 3.475 4.55 2.825 3.025 2.375 2.175 Acc (%) [75] $\uparrow$ 0.0 0.0 30.0 22.5 27.5 22.5 30.0 32.5 ImageReward [77] $\uparrow$ 0.680 0.526 0.954 0.889 0.803 0.789 0.840 0.845 VQA [78] $\uparrow$ 0.920 0.859 0.934 0.920 0.916 0.928 0.922 0.930 Aesthetic Preference $\text{Aesthetic}$$^{\dagger}$ [76] $\uparrow$ 6.432 6.281 7.436 7.324 7.452 7.343 7.412 7.469 ImageReward [77] $\uparrow$ 1.106 1.045 1.233 1.258 1.144 1.367 1.217 1.262 VQA [78] $\uparrow$ 0.891 0.902 0.894 0.907 0.888 0.905 0.904 0.909
@@
FreeDoM [40] TDS [14] DAS [12] Top-$K$-of-NN ULA MALA $\Psi$-Sampler
@@
Appendix I Additional Qualitative Results
@@
Layout FreeDoM [40] TDS [14] DAS [12] Top-$K$-of-NN ULA MALA $\Psi$-Sampler
@@
FreeDoM [40] TDS [14] DAS [12] Top-$K$-of-NN ULA MALA $\Psi$-Sampler
@@
FreeDoM [40] TDS [14] DAS [12] Top-$K$-of-NN ULA MALA $\Psi$-Sampler
@@
FreeDoM [40] TDS [14] DAS [12] Top-$K$-of-NN ULA MALA $\Psi$-Sampler